OpenStudy (anonymous):

When 2x+3y=6 is in slope-intercept form, the y-intercept is...?

OpenStudy (ipwnbunnies):

Maybe you should first get the equation in slope-intercept form...

OpenStudy (anonymous):

isolate for y so put in y=mx+b form.

OpenStudy (anonymous):

2x-3y=6 (-2x)+2x-3y=6+(-2x) -3y=-2x+6 -3y(1/-3)= -2x(1/-3)+6(1/-3) y= -2/3x-2

OpenStudy (anonymous):

y intercept is -2

OpenStudy (anonymous):

Alright thanks alot guys @Bkfrostyblue @CarmenK